# A $2$-distance $(2\Delta+7)$-coloring of planar graphs

### v2

@@ Abstract.
A vertex coloring of a graph $G$ is called a $2$-distance coloring if any two vertices at a distance at most $2$ from each other receive different colors. Recently, Bousquet et al. (Discrete Mathematics, 346(4), 113288, 2023) proved that $72\Delta+$ 7 colors are sufficient for the $2$-distance coloring of planar graphs with maximum degree $\Delta\geq 9$. In this paper, we strengthen their result by removing the maximum degree constraint and show that all planar graphs admit a 2-distance $(2\Delta+7)$-coloring. This particularly improves the result of Van den Heuvel and McGuinness (Journal of Graph Theory, 42(2), 110-124, 2003).
@@ 1. Introduction
Among the numerous problems and concepts associated with graphs, coloring stands out as a fundamental and extensively studied topic. A vertex coloring of a graph involves assigning colors to its vertices to ensure that adjacent vertices have distinct colors. One particularly interesting variant of vertex coloring is 2-distance coloring, where any two vertices at a distance of two have distinct colors. Its motivation arises from the necessity in some real-life problems to assign different colors not only to adjacent vertices but also to those in close proximity [11, 14]. The concept of 2-distance coloring was first introduced in [9, 10], and it has received considerable attention due to the fact that certain problems, such as the Total Coloring Conjecture (see [4] for details), can be formulated as a 2-distance coloring of specific graphs. A comprehensive survey on 2-distance coloring and its related coloring concepts has been presented by Cranston [4].
@@
Wegner’s Conjecture is one of the most popular problems in graph coloring, and it has remained open for many years, except for the case when $\Delta=3$, which was solved by Thomassen [16] (independently by Hartke et al. [7]). For general planar graphs, Van den Heuvel and McGuinness [8] showed that $\chi_{2}(G)\leq 2\Delta+25$, while the bound $\chi_{2}(G)\leq\big{\lceil}\frac{5\Delta}{3}\big{\rceil}+78$ was proved by Molloy and Salavatipour [15]. On the other hand, some improved results are presented in [2, 3, 5, 6, 13] with a certain degree or girth restrictions. In particular, Bousquet et al. [1] recently showed that $\chi_{2}(G)\leq 2\Delta+7$ if $\Delta\geq 9$, which significantly improves upon the work in [12].
In this paper, we strengthen the result of Bousquet et al. [1] by removing the maximum degree constraint and show that all planar graphs admit a 2-distance $(2\Delta+7)$-coloring, which particularly improves the best-known bound of $252\Delta+$ 25 provided by Van den Heuvel and McGuinness [8].
@@
All graphs in this paper are assumed to be simple, and we refer to [18] for terminology and notation not defined here. When $G$ is a graph, we use $V(G),E(G),F(G),$ and $\Delta(G)$ to denote the vertex, edge and face set, and the maximum degree of $G$, respectively. If there is no confusion in the context, we abbreviate $\Delta(G)$ to $\Delta$. For a positive integer $t$, we denote by $[t]$ the set of integers between $1$ and $t$. Given a planar graph $G$, we denote by $\ell(f)$ the length of a face $f$ and by $d(v)$ the degree of a vertex $v$. A $k$-vertex is a vertex of degree $k$. A $k^{-}$-vertex is a vertex of degree at most $k$ while a $k^{+}$-vertex is a vertex of degree at least $k$. A $k$ ($k^{-}$ or $k^{+}$)-face is defined similarly. A vertex $u\in N(v)$ is called $k$-neighbour (resp. $k^{-}$-neighbour, $k^{+}$-neighbour) of $v$ if $d(u)=k$ (resp. $d(u)\leq k$, $d(u)\geq k$).
[…]
@@
Let $G$ be a minimal counterexample to Theorem 1.2 such that $|V(G)|+|E(G)|$ is minimum. Thus, $G$ does not admit any $2$-distance $(2\Delta+7)$-coloring, but any planar graph $G^{\prime}$ obtained from $G$ with smaller $|V(G^{\prime})|+|E(G^{\prime})|$ admits a 2-distance $(2\Delta+7)$-coloring.
Clearly, $G$ is a connected graph. Due to the results of Bousquet et al. [1] and Zou et al. [19] (also [5]), we only need to consider the case $6\leq\Delta\leq 8$.
We call a graph $H$ proper with respect to $G$ if $H$ is obtained from $G$ by deleting some edges or vertices and adding some edges, ensuring that for every pair of vertices $x_{1}$ and $x_{2}$ in $V(G)\cap V(H)$ having distance at most 2 in $G$, they also have distance at most $2$ in $H$. If $f$ is a 2-distance coloring of such a graph $H$, then $f$ can be extended to the whole graph $G$, provided that each of the remaining uncolored vertices has an available color.
For a given vertex $v$ with $d(v)=k$, we assume that $v_{1},v_{2},\ldots,v_{k}$ are the neighbours of $v$ in clockwise order throughout this section.
@@ 2.1. The case $\Delta=6$
Recall that $G$ does not admit a $2$-distance $19$-coloring, but any planar graph $G^{\prime}$ obtained from $G$ with a smaller $|V(G^{\prime})|+|E(G^{\prime})|$ admits a 2-distance $19$-coloring.
@@
If $v$ is a vertex of degree at most $2$, then $d_{2}(v)\leq 12$. Let $G^{\prime}$ be the graph obtained from $G-v$ by adding an edge between the vertices in $N(v)$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring $f$ with $19$ colors. Notice that $G^{\prime}$ is proper with respect to $G$. Since $d_{2}(v)\leq 12$, we can color $v$ with an available color, a contradiction. ∎
@@
Let $v$ be a $3$-vertex. Then $m_{3}(v)=0$, $m_{4}(v)\leq 1$, and each neighbour of $v$ is a $6(4^{-})$-vertex.
@@
Let $v$ be a $3$-vertex, and denote by $v_{1},v_{2},v_{3}$ the neighbours of $v$ in clockwise order. Assume for a contradiction that $v$ is incident to a $3$-face, say $f_{1}=v_{1}vv_{2}$. It is clear that $d_{2}(v)\leq 16$. If we set $G^{\prime}=G-v+\{v_{1}v_{3}\}$, then $G^{\prime}$ would be proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring with $19$ colors. Since $d_{2}(v)\leq 16$, we can color $v$ with an available color, a contradiction.
Suppose now that $m_{4}(v)\geq 2$, i.e., $v$ is incident to two $4$-faces, say $f_{1}=v_{1}vv_{2}x$ and $f_{2}=v_{2}vv_{3}y$. Clearly, we have $d_{2}(v)\leq 16$. If we set $G^{\prime}=G-v+\{v_{1}v_{3}\}$, then $G^{\prime}$ would be proper with respect to $G$. Similarly as above, we get a contradiction.
Next we claim that each neighbour of $v$ is a $6(4^{-})$-vertex. Note that $v$ cannot have any $6(5^{+})$-neighbour since $m_{3}(v)=0$. Moreover, $v$ cannot have any $5^{-}$-neighbour. Indeed, if $v_{1}$ is such a vertex, then we set $G^{\prime}=G-v+\{v_{1}v_{2},v_{1}v_{3}\}$. In such a case, $d_{2}(v)\leq 17$, and $G^{\prime}$ is proper with respect to $G$. Similarly as above, we get a contradiction. Thus, each neighbour of $v$ is a $6(4^{-})$-vertex. ∎
We now address some vertices having some special forms: If $v$ is a $4(1)$- or a $4(2)$-vertex, then we call $v$ as a bad $4$-vertex. Similarly, if $v$ is a $5(4)$- or a 5(5)555(5)5 ( 5 )-vertex, then we call $v$ as a bad $5$-vertex.
@@
If $v$ is a $4$-vertex, then $m_{3}(v)\leq 2$, and $v$ has no 6(6)666(6)6 ( 6 )-neighbour.
@@
Let $v$ be a $4$-vertex, and denote by $f_{1},f_{2},\ldots,f_{4}$ the faces incident to $v$. Assume for a contradiction that $m_{3}(v)\geq 3$. It follows that we have $f_{i}=v_{i}vv_{i+1}$ for $i\in[3]$. Clearly, $d_{2}(v)\leq 18$. If we set $G^{\prime}=G-v+\{v_{1}v_{4}\}$ (assuming $v_{1}v_{4}\notin E(G)$), then, by minimality, the graph $G^{\prime}$ has a 2-distance coloring with $19$ colors. Observe that every vertex in $V(G)$, except $v$, is colored with the same color that appears on the same vertex in $G^{\prime}$. Namely, $G^{\prime}$ is proper with respect to $G$. Since $d_{2}(v)\leq 18$, we assign $v$ with an available color, a contradiction.
Suppose now that $v$ has a 6(6)666(6)6 ( 6 )-neighbour. Without loss of generality, assume that $v_{2}$ is a 6(6)666(6)6 ( 6 )-vertex. In such a case, $v$ must be incident to two adjacent $3$-faces $f_{1},f_{2}$, say $f_{1}=v_{1}vv_{2}$ and $f_{2}=v_{2}vv_{3}$. Moreover, both $v_{1}v_{2}$ and $v_{2}v_{3}$ must be contained in two $3$-faces. This implies that $d_{2}(v)\leq 18$. If we set $G^{\prime}=G-v+\{v_{2}v_{4}\}$, then $G^{\prime}$ would be proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring with $19$ colors. Since $d_{2}(v)\leq 18$, we assign $v$ with an available color, a contradiction. ∎
@@
Let $v$ be a $4(1)$-vertex. If $v$ has a $4$-neighbour, then the other neighbours of $v$ are different from $4$- and $5(4)$-vertices.
@@
Suppose that $v$ has a $4$-neighbour, say $v_{1}$, and assume by contradiction that $v$ has another neighbour consisting of $4$- or $5(4)$-vertices. In this case, we have $d_{2}(v)\leq 18$. If we set $G^{\prime}=G-v+\{v_{1}v_{2},v_{1}v_{3},v_{1}v_{4}\}$, then $G^{\prime}$ would be proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring with $19$ colors. Since $d_{2}(v)\leq 18$, we can color $v$ with an available color, a contradiction. ∎
@@
Let $v$ be a $4(1)$-vertex, and denote by $f_{1},f_{2},\ldots,f_{4}$ the faces incident to $v$. Suppose that $f_{1}$ is a $3$-face with $f_{1}=v_{1}vv_{2}$. Let $m_{4}(v)=r$ for $0\leq r\leq 3$. Assume for a contradiction that $n_{6}(v)\leq r$. We then deduce that $d_{2}(v)\leq 18$, since $v$ has $4-r$ many $5^{-}$-neighbours, and $v$ is incident to $r$ many $4$-faces. This particularly implies that either $v$ has two $5^{-}$-neighbours or $v$ is incident to three $4$-faces.
If $v$ has two $5^{-}$-neighbours $v_{i},v_{j}$ such that $\{i,j\}\cap\{1,2\}\neq\emptyset$, then we set $G^{\prime}=G-v+\{v_{k}v_{3},v_{k}v_{4}\}$ for $k\in\{i,j\}\cap\{1,2\}$. If $v$ has two $5^{-}$-neighbours $v_{i},v_{j}$ with $i<j$ such that $\{i,j\}\cap\{1,2\}=\emptyset$, then we set $G^{\prime}=G-v+\{v_{i}v_{j},v_{i}v_{2},v_{j}v_{1}\}$. If $v$ is incident to exactly three $4$-faces, then we set $G^{\prime}=G-v+\{v_{2}v_{3},v_{1}v_{4}\}$. In each case, we observe that $G^{\prime}$ is proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring with $19$ colors. Since $d_{2}(v)\leq 18$, we can color $v$ with an available color, a contradiction. ∎
@@
Let $v$ be a $4(2)$-vertex, and denote by $f_{1},f_{2},\ldots,f_{4}$ the faces incident to $v$. Suppose that $f_{1}$ and $f_{2}$ are $3$-faces such that $f_{1}=v_{i}vv_{i+1}$ and $f_{2}=v_{j}vv_{j+1}$ with $i<j$ and $i,j\in[4]$ in a cyclic fashion. Assume for a contradiction that $v$ is incident to two $4$-faces. In this case, we have $d_{2}(v)\leq 18$. If $f_{1},f_{2}$ are adjacent, then we set $G^{\prime}=G-v+\{v_{j}v_{k}\}$ for $k\in N(v)\setminus\{v_{i},v_{j},v_{j+1}\}$. Else, we set $G^{\prime}=G-v+\{v_{i}v_{j+1},v_{i+1}v_{j}\}$. In both cases, $G^{\prime}$ is proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring with $19$ colors. Since $d_{2}(v)\leq 18$, we can color $v$ with an available color, a contradiction.
On the other hand, if $m_{4}(v)=0$ and $n_{6}(v)\leq 2$, then $v$ has two $5^{-}$-neighbours, and so $d_{2}(v)\leq 18$. Similarly, if $m_{4}(v)=1$ and $n_{6}(v)\leq 3$, then $v$ has a $5^{-}$-neighbour, and so $d_{2}(v)\leq 18$, as $m_{4}(v)=1$. By applying the same argument as above, we get a contradiction. ∎
@@
Let $v$ be a $4$-vertex with $m_{3}(v)=2$. Suppose first that one of $v_{i}$’s is a $4$-vertex, say $v_{1}$. Obviously, we have $d_{2}(v)\leq 18$, and let $G^{\prime}=G-v+\{v_{1}v_{2},v_{1}v_{3},v_{1}v_{4}\}$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring $f$ with $19$ colors. Observe that $G^{\prime}$ is proper with respect to $G$. Since $d_{2}(v)\leq 18$, we can color $v$ with an available color, a contradiction.
Suppose now that $v$ has a bad $5$-neighbour, say $v_{i}$. Then, there exists an edge $v_{i}v_{j}$ for $v_{j}\in N(v)$ such that $v_{i}v_{j}$ is contained in two $3$-faces. It follows that we have $d_{2}(v)\leq 18$. If we set $G^{\prime}=G-v+\{v_{i}v_{p},v_{i}v_{t}\}$ for $v_{p},v_{t}\in N(v)\setminus\{v_{i},v_{j}\}$, then $G^{\prime}$ would be proper with respect to $G$. Similarly as above, we get a contradiction. ∎
@@
Let $v$ be a $5(4)$-vertex. If $v$ has a $4$-neighbour, then the other neighbours of $v$ are different from $4$- and $5(4)$-vertices.
@@
Let $v$ be a $5(4)$-vertex and denote by $f_{1},f_{2},\ldots,f_{5}$ the faces incident to $v$ such that $f_{i}=v_{i}vv_{i+1}$ for $i\in[4]$. Suppose that $v$ has a $4$-neighbour, and assume by contradiction that $v$ has another neighbour consisting of $4$- or $5(4)$-vertices. We then deduce that $d_{2}(v)\leq 18$. If we set $G^{\prime}=G-v+\{v_{1}v_{5}\}$, then $G^{\prime}$ would be proper with respect to $G$. By minimality, the graph $G^{\prime}$ has a 2-distance coloring with $19$ colors. Since $d_{2}(v)\leq 18$, we can color $v$ with an available color, a contradiction. ∎
The following is a direct consequence of Lemma 2.8, since a $5(4)$-vertex cannot have both a $4$-neighbour and a $5(4)$-neighbour simultaneously.
@@
    If $m_{4}(v)=1$, then $v$ has three $6(5^{-})$-neighbours.

### v3

@@ Abstract.
A vertex coloring of a graph $G$ is called a $2$-distance coloring if any two vertices at a distance at most $2$ from each other receive different colors. Recently, Bousquet et al. (Discrete Mathematics, 346(4), 113288, 2023) proved that $2\Delta+7$ colors are sufficient for the $2$-distance coloring of planar graphs with maximum degree $\Delta\geq 9$. In this paper, we strengthen their result by removing the maximum degree constraint and show that all planar graphs admit a 2-distance $(2\Delta+7)$-coloring. This particularly improves the result of Van den Heuvel and McGuinness (Journal of Graph Theory, 42(2), 110-124, 2003).
@@ 1. Introduction
Among the numerous problems and concepts associated with graphs, coloring stands out as a fundamental and extensively studied topic. A vertex coloring of a graph assigns colors to its vertices so that adjacent vertices receive distinct colors. One particularly interesting variant of vertex coloring is 2-distance coloring, where any two vertices at a distance of $2$ or less have distinct colors. Its motivation arises from the necessity in some real-life problems to assign different colors not only to adjacent vertices but also to those in close proximity [14, 16]. The concept of 2-distance coloring was first introduced in [12, 13], and it has received considerable attention due to the fact that certain problems, such as the Total Coloring Conjecture (see [5] for details), can be formulated as a 2-distance coloring of specific graphs. A comprehensive survey on 2-distance coloring and its related coloring concepts has been presented by Cranston [5].
@@
Wegner’s Conjecture is one of the most popular problems in graph coloring, and it has remained open for many years, except for the case when $\Delta=3$, which was solved by Thomassen [18] (independently by Hartke et al. [9]). The conjecture is also known to be asymptotically true for large values of $\Delta$, as shown in [1, 10]. For general planar graphs, Van den Heuvel and McGuinness [11] showed that $\chi_{2}(G)\leq 2\Delta+25$, while the bound $\chi_{2}(G)\leq\big\lceil\frac{5\Delta}{3}\big\rceil+78$ was proved by Molloy and Salavatipour [17]. On the other hand, some improved results are presented in [3, 4, 6, 7, 8] with a certain degree or girth restrictions. In particular, Bousquet et al. [2] recently showed that $\chi_{2}(G)\leq 2\Delta+7$ if $\Delta\geq 9$, which significantly improves upon the work in [15].
In this paper, we strengthen the result of Bousquet et al. [2] by removing the maximum degree constraint and show that all planar graphs admit a 2-distance $(2\Delta+7)$-coloring, which particularly improves the best-known bound of $2\Delta+25$ provided by Van den Heuvel and McGuinness [11].
@@
All graphs in this paper are assumed to be simple, and we refer to [20] for terminology and notation not defined here. When $G$ is a graph, we use $V(G),E(G),F(G),$ and $\Delta(G)$ to denote the vertex, edge and face set, and the maximum degree of $G$, respectively. If there is no confusion in the context, we abbreviate $\Delta(G)$ to $\Delta$. For a positive integer $t$, we denote by $[t]$ the set of integers between $1$ and $t$. Given a planar graph $G$, we denote by $\ell(f)$ the length of a face $f$ and by $d(v)$ the degree of a vertex $v$. A $k$-vertex is a vertex of degree $k$. A $k^{-}$-vertex is a vertex of degree at most $k$ while a $k^{+}$-vertex is a vertex of degree at least $k$. A $k$ ($k^{-}$ or $k^{+}$)-face is defined analogously as for the vertices, where the degree of a face is the number of edges incident to it. A vertex $u\in N(v)$ is called $k$-neighbour (resp. $k^{-}$-neighbour, $k^{+}$-neighbour) of $v$ if $d(u)=k$ (resp. $d(u)\leq k$, $d(u)\geq k$).
For a vertex $v\in V(G)$, we use $n_{i}(v)$ to denote the number of $i$-vertices adjacent to $v$. We denote by $d(u,v)$ the distance between $u$ and $v$ for a pair $u,v\in V(G)$. Also, we set $N_{i}(v):=\{u\in V(G)\ |\ 1\leq d(u,v)\leq i\}$ for $i\geq 1$, so $N_{1}(v)=N(v)$ and let $d_{2}(v)=|N_{2}(v)|$. For $v\in V(G)$, we use $m_{k}(v)$ to denote the number of $k$-faces incident with $v$. A $k$-vertex $v$ with $m_{3}(v)=d$ is called $k(d)$-vertex. In particular, a $k$-vertex $v$ is called $k(d^{-})$-vertex (resp. $k(d^{+})$-vertex) if $m_{3}(v)\leq d$ (resp. $m_{3}(v)\geq d$). Two faces $f_{1}$ and $f_{2}$ are said to be adjacent if they share a common edge.
@@
Let $G$ be a minimal counterexample to Theorem 1.2 such that $|V(G)|+|E(G)|$ is minimum. Thus, $G$ does not admit any $2$-distance $(2\Delta+7)$-coloring, whereas every planar graph $G^{\prime}$ obtained from $G$ with $|V(G^{\prime})|+|E(G^{\prime})|<|V(G)|+|E(G)|$ admits a $2$-distance $(2\Delta+7)$-coloring. By the results of [2, 6], it suffices to consider the case $6\leq\Delta\leq 8$. Clearly, $G$ is connected.
We call a graph $H$ proper with respect to $G$ if $H$ is obtained from $G$ by deleting some edges or vertices and adding some edges, in such a way that every pair of vertices $x_{1},x_{2}\in V(G)\cap V(H)$ that have distance at most $2$ in $G$ also have distance at most $2$ in $H$. If $f$ is a $2$-distance coloring of such a graph $H$, then $f$ can be extended to a $2$-distance coloring of $G$, provided that each remaining uncolored vertex has an available color. By the minimality of $G$, we obtain the following.
Remark 2.1.
If $H$ is proper with respect to $G$ such that $V(G)\setminus V(H)=\{v\}$, then $d_{2}(v)\geq 2\Delta+7$.
Proof.
Assume, for a contradiction, that there exists a vertex $v\in V(G)\setminus V(H)$ such that $d_{2}(v)\leq 2\Delta+6$. Since $H$ is proper with respect to $G$, by the minimality of $G$, the graph $H$ admits a $2$-distance coloring using $2\Delta+7$ colors. As $d_{2}(v)\leq 2\Delta+6$, there is at least one available color for $v$, and thus the coloring of $H$ extends to a $2$-distance coloring of $G$, a contradiction to the minimality of $G$. ∎
For a vertex $v$ with $d(v)=k$, let $v_{1},v_{2},\ldots,v_{k}$ denote the neighbors of $v$ in clockwise order, and let $f_{1},f_{2},\ldots,f_{k}$ be the faces incident to $v$. Define $E(v)=\{v_{i}v_{i+1}\in E(G):i\in[k]\,\}$, where indices are taken modulo $k$. Let $t(v)$ denote the number of edges in $E(v)$ that are contained in two $3$-faces. We begin with a useful lemma that provides an upper bound on $d_{2}(v)$.
Lemma 2.2.
For any vertex $v$, we have $d_{2}(v)\leq\left(\sum_{u\in N(v)}d(u)\right)-2m_{3}(v)-m_{4}(v)-t(v)$.
Proof.
Recall that $d_{2}(v)$ counts the number of vertices at distance at most $2$ from $v$, that is, $d_{2}(v)=|N_{2}(v)|$. Let $p=\sum_{u\in N(v)}d(u)$. From $p$, we subtract $2$ for each $3$-face incident to $v$ and $1$ for each $4$-face incident to $v$. Furthermore, for each edge $v_{i}v_{i+1}$, if $v_{i}$ and $v_{i+1}$ have a common neighbor other than $v$, we subtract an additional $1$ from $p$. Consequently, we obtain $d_{2}(v)\leq p-2m_{3}(v)-m_{4}(v)-t(v).$
In the remainder of the paper, we divide the proof into three cases according to $\Delta\in\{6,7,8\}$. In each case, we employ the same overall strategy: we first derive structural properties from the minimality of $G$, and then introduce an appropriate set of discharging rules and apply the discharging method to reach a contradiction, thereby showing that $G$ cannot exist.
@@ 2.1. The case $\Delta=6$
Recall that $G$ does not admit a $2$-distance 1919-coloring, whereas every planar graph $G^{\prime}$ obtained from $G$ with a smaller value of $|V(G^{\prime})|+|E(G^{\prime})|$ admits one.
@@
If $v$ is a vertex of degree at most $2$, then $d_{2}(v)\leq 12$. Let $G^{\prime}$ be the graph obtained from GvG-v by adding an edge between the vertices in $N(v)$. Notice that $G^{\prime}$ is proper with respect to $G$. By Remark 2.1, this yields a contradiction. ∎
@@
Let $v$ be a $3$-vertex. Then $m_{3}(v)=0$, $m_{4}(v)\leq 1$, and every neighbor of $v$ is a $6(4^{-})$-vertex.
@@
Assume first, for a contradiction, that $v$ is incident to a $3$-face, say $f_{1}=v_{1}vv_{2}$. By Lemma 2.2, we have $d_{2}(v)\leq 16$. If we set $G^{\prime}=G-v+\{v_{1}v_{3}\}$, then $G^{\prime}$ is proper with respect to $G$, yielding a contradiction to Remark 2.1.
Suppose next that $m_{4}(v)\geq 2$, that is, $v$ is incident to two $4$-faces, say $f_{1}=v_{1}vv_{2}x$ and $f_{2}=v_{2}vv_{3}y$. Again, $d_{2}(v)\leq 16$. If we set $G^{\prime}=G-v+\{v_{1}v_{3}\}$, then $G^{\prime}$ is proper with respect to $G$, giving the same contradiction as above.
We now show that every neighbor of $v$ is a $6(4^{-})$-vertex. Since $m_{3}(v)=0$, the vertex $v$ cannot have a $6(5^{+})$-neighbor. Moreover, $v$ cannot have a $5^{-}$-neighbor. Indeed, suppose $v_{1}$ is such a neighbor. Let $G^{\prime}=G-v+\{v_{1}v_{2},v_{1}v_{3}\}$. In this case $d_{2}(v)\leq 17$, and $G^{\prime}$ is proper with respect to $G$, again contradicting Remark 2.1. Hence every neighbor of $v$ must be a $6(4^{-})$-vertex. ∎
We now introduce some terminology for vertices of special types. A $4(1)$- or $4(2)$-vertex is called a bad $4$-vertex. Similarly, a $5(4)$- or $5(5)$-vertex is called a bad $5$-vertex.
@@
If $v$ is a $4$-vertex, then $m_{3}(v)\leq 2$, and $v$ has no $k(k)$-neighbour for any $k\leq 6$.
@@
Let $v$ be a $4$-vertex. Assume, for a contradiction, that $m_{3}(v)\geq 3$. Let $f_{i}=v_{i}vv_{i+1}$ for $i\in[3]$. By Lemma 2.2, we have $d_{2}(v)\leq 18$. If we define $G^{\prime}=G-v+\{v_{1}v_{4}\}$ (assuming $v_{1}v_{4}\notin E(G)$), then $G^{\prime}$ is proper with respect to $G$. This contradicts Remark 2.1.
Now suppose that $v$ has a $k(k)$-neighbour for some $k\leq 6$. Without loss of generality, assume that $v_{2}$ is a $k(k)$-vertex. Then $v$ must be incident with two consecutive $3$-faces $f_{1}$ and $f_{2}$, say $f_{1}=v_{1}vv_{2}$ and $f_{2}=v_{2}vv_{3}$. Moreover, each of the edges $v_{1}v_{2}$ and $v_{2}v_{3}$ must be contained in two $3$-faces. Hence $d_{2}(v)\leq 18$ by Lemma 2.2. If we set $G^{\prime}=G-v+\{v_{2}v_{4}\}$, then $G^{\prime}$ would again be proper with respect to $G$, giving the same contradiction. ∎
@@
Let $v$ be a $4(1)$-vertex. If $v$ has a $4$-neighbour, then the other neighbours of $v$ are neither $4$-vertices nor bad $5$-vertices.
@@
Suppose that $v$ has a $4$-neighbour, say $v_{1}$, and assume for a contradiction that $v$ has another neighbour that is either a $4$-vertex or a $5(4^{+})$-vertex. In this case, Lemma 2.2 implies $d_{2}(v)\leq 18$. If we form $G^{\prime}=G-v+\{v_{1}v_{2},\,v_{1}v_{3},\,v_{1}v_{4}\}$, then $G^{\prime}$ is proper with respect to $G$. By Remark 2.1, this yields a contradiction. ∎
@@
Let $f_{1}=v_{1}vv_{2}$ be the unique $3$-face incident to $v$. Assume $m_{4}(v)=r$ for some $0\leq r\leq 3$. Suppose, for a contradiction, that $n_{6}(v)\leq r$. Then $v$ has 4r4-r neighbours of degree at most $5$, and is incident to $r$ 44-faces. By Lemma 2.2, this implies $d_{2}(v)\leq 18$. In particular, either $v$ has two $5^{-}$-neighbours, or $v$ is incident to three $4$-faces. We show that in each case, we can construct a graph $G^{\prime}$ proper with respect to $G$, which gives a contradiction.
If $v$ has two $5^{-}$-neighbours $v_{i},v_{j}$ such that $\{i,j\}\cap\{1,2\}\neq\emptyset$, then we set $G^{\prime}=G-v+\{v_{k}v_{3},v_{k}v_{4}\}$ for $k\in\{i,j\}\cap\{1,2\}$. If $v$ has two $5^{-}$-neighbours $v_{i},v_{j}$ with $i<j$ such that $\{i,j\}\cap\{1,2\}=\emptyset$, then we set $G^{\prime}=G-v+\{v_{i}v_{j},v_{i}v_{2},v_{j}v_{1}\}$. If $v$ is incident to exactly three $4$-faces, then we set $G^{\prime}=G-v+\{v_{2}v_{3},v_{1}v_{4}\}$. In each case, $G^{\prime}$ is proper with respect to $G$. By Remark 2.1, this yields a contradiction. Hence, $n_{6}(v)\geq r+1$, as desired. ∎
@@
Suppose that $f_{1}$ and $f_{2}$ are $3$-faces incident to $v$, where $f_{1}=v_{i}vv_{i+1}$ and $f_{2}=v_{j}vv_{j+1}$ with $i<j$, and $i,j\in[4]$ taken cyclically. Assume, for a contradiction, that $v$ is incident to two $4$-faces. Then, by Lemma 2.2, we have $d_{2}(v)\leq 18$. If $f_{1}$ and $f_{2}$ are adjacent, choose $v_{k}\in N(v)\setminus\{v_{i},v_{i+1},v_{j},v_{j+1}\}$ and set $G^{\prime}=G-v+\{v_{j}v_{k}\}$. Otherwise, set $G^{\prime}=G-v+\{v_{i}v_{j+1},\,v_{i+1}v_{j}\}$. In both cases, the resulting graph $G^{\prime}$ is proper with respect to $G$. By Remark 2.1, this yields a contradiction. Thus $m_{4}(v)\leq 1$.
Now assume $m_{4}(v)=0$ and $n_{6}(v)\leq 2$. Then $v$ has two $5^{-}$-neighbours, and so Lemma 2.2 gives $d_{2}(v)\leq 18$. Similarly, if $m_{4}(v)=1$ and $n_{6}(v)\leq 3$, then $v$ has a $5^{-}$-neighbour, and again $d_{2}(v)\leq 18$. In each case, applying the same modification of $G$ as above yields a proper graph $G^{\prime}$, contradicting Remark 2.1. Thus $n_{6}(v)\geq r+3$ for $r=m_{4}(v)$. ∎
@@
Suppose first that one of $v_{i}$’s is a $4$-vertex, say $v_{1}$. Obviously, we have $d_{2}(v)\leq 18$, and let $G^{\prime}=G-v+\{v_{1}v_{2},v_{1}v_{3},v_{1}v_{4}\}$, assuming these edges are not already present. Observe that $G^{\prime}$ is proper with respect to $G$. This contradicts Remark 2.1.
Suppose now that $v$ has a bad $5$-neighbour, say $v_{i}$. Then, there exists an edge $v_{i}v_{j}$ for $v_{j}\in N(v)$ such that $v_{i}v_{j}$ is contained in two $3$-faces. By Lemma 2.2, this implies $d_{2}(v)\leq 18$. If we set $G^{\prime}=G-v+\{v_{i}v_{p},v_{i}v_{t}\}$ for $v_{p},v_{t}\in N(v)\setminus\{v_{i},v_{j}\}$, then $G^{\prime}$ would be proper with respect to $G$. Similarly as above, we get a contradiction. ∎
@@
Let $v$ be a $5(4)$-vertex. If $v$ has a $4$-neighbour, then the other neighbours of $v$ are neither $4$-vertices nor $5(4)$-vertices.
@@
Let $f_{i}=v_{i}vv_{i+1}$ for $i\in[4]$. Suppose that $v$ has a $4$-neighbour, and assume for a contradiction that $v$ has another neighbour that is either a $4$-vertex or a $5(4)$-vertex. Then, by Lemma 2.2, we obtain $d_{2}(v)\leq 18$. If we define $G^{\prime}=G-v+\{v_{1}v_{5}\}$, then $G^{\prime}$ is proper with respect to $G$. By Remark 2.1, this yields a contradiction. ∎
The following is an immediate consequence of Lemma 2.10, since a $5(4)$-vertex cannot simultaneously have a $4$-neighbour and a $5(4)$-neighbour.
@@
    If $m_{4}(v)=1$, then $v$ has at least three $6(5^{-})$-neighbours.